OpenStudy (marinadanyel):

What is the discontinuity and zero of the function f(x) = the quantity of 2 x squared plus 5 x minus 12, all over x plus 4 ? Discontinuity at (−4, −11), zero at ( three halves , 0) Discontinuity at (−4, −11), zero at ( negative three halves , 0) Discontinuity at (4, 5), zero at ( three halves , 0) Discontinuity at (4, 5), zero at ( negative three halves , 0)

TheSmartOne (thesmartone):

the x-coordinate of the discontinuity can be found by setting the denominator equal to 0

OpenStudy (marinadanyel):

so, x = - 4?

TheSmartOne (thesmartone):

yes, and can you write the equation using the equation tool or the draw tool so I can see it better?

OpenStudy (marinadanyel):

@thesmartone

TheSmartOne (thesmartone):

can you factor the numerator?

OpenStudy (marinadanyel):

Yes, (2x - 3) • (x + 4).

TheSmartOne (thesmartone):

now set each equal to 0 2x - 3 = 0 x +4 = 0 and those are the roots

TheSmartOne (thesmartone):

roots are the same thing as zeroes or solutions or x-intercepts :)

OpenStudy (marinadanyel):

Oh, okay, so where do they come in for the answer?

OpenStudy (marinadanyel):

I don't understand how to get the answer. :(

TheSmartOne (thesmartone):

2x - 3 = 0 what is x = ?

OpenStudy (marinadanyel):

0?

TheSmartOne (thesmartone):

no, add 3 to both sides and then divide by 2

OpenStudy (marinadanyel):

Oh! \[\frac{ 3 }{ 2 }\]

OpenStudy (marinadanyel):

So my answer is A. Discontinuity at (−4, −11), zero at ( three halves , 0) ???

OpenStudy (marinadanyel):

@TheSmartOne Is that correct?

TheSmartOne (thesmartone):

correct! ^-^

OpenStudy (marinadanyel):

Thank you! (^̮^)